# A Survey of Controllable Learning: Methods and Applications in Information Retrieval

### v2

@@ 2 Foundations of Controllable Learning
In this section, we first give a formal definition of controllable learning in terms of task requirements, and describe the procedure of controllable learning.
@@
The concept of controllable learning is manifested across various AI research directions, yet there is currently lack of a unified formal definition. To explicitly express the ingredients of controllable learning and unify existing controllable learning methods, we first define controllable learning in terms of task specifications. In brief, controllable learning is the ability to find a learner that can adapt to different task requirements without the need for retraining, thereby meeting the desired task targets, as detailed in Definition 2.1.
Definition 2.1 (Controllable Learning).
@@
Define a task requirement triplet $\mathcal{T}=\{\bm{s}_{\mathrm{desc}},\bm{s}_{\mathrm{ctx}},\bm{s}_{\mathrm{tgt}}\}\in\Gamma$, where $\bm{s}_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ represents the task description, $\bm{s}_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ represents the context related to the task, and $\bm{s}_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ represents the task target. Given an input space $\mathcal{X}$ and an output space $\mathcal{Y}$, for a learner $f:\mathcal{X}\rightarrow\mathcal{Y}$, controllable learning aims to find a control function $h$ that maps the learner $f$, the task description $\bm{s}_{\mathrm{desc}}\in\mathcal{T}$, and the context $\bm{s}_{\mathrm{ctx}}\in\mathcal{T}$ to a new learner $f_{\mathcal{T}}$ that fulfills the task target $\bm{s}_{\mathrm{tgt}}\in\mathcal{T}$, i.e,
@@
Unlike domain adaptation or transfer learning[69, 70, 71], controllable learning adjusts the model adaptively in the face of new task requirements, eliminating the need for retraining for new tasks during the deployment phase. More specifically, in the context of a controllable learner during the testing phase, adjustments based on the task description can be categorized as follows: altering model inputs qualifies as a pre-processing method, modifying model parameters constitutes an in-processing method, and changing model outputs can be regarded as a post-processing method. However, it also presents challenges during the training phase in controllable learning, particularly in guiding the control function to identify and adapt to various task requirements effectively.
For the task requirement in Definition 2.1, the task target $\bm{s}_{\mathrm{tgt}}$ can be considered as the ideal quantitative metric that the controllable learner aims to achieve. The task description $\bm{s}_{\mathrm{desc}}$ is the specific representation of $\bm{s}_{\mathrm{tgt}}$ that can be inputted into an algorithm. $\bm{s}_{\mathrm{desc}}$ can be represented as a natural language prompt or as a vector containing multiple weights, among other forms. The context $\bm{s}_{\mathrm{ctx}}$ in task requirements can consist of features like historical data and the current user’s profile, offering controllable learners more background knowledge for prediction.
@@
Under the definition, the procedure of controllable learning could be shown as Figure 2. When the output space $\mathcal{Y}$ in Definition 2.1 is not labels or lists but instead generated content, it is referred to as controllable generation. In this survey, our primary focus lies on controllable learning methods within information retrieval applications, which typically belong to controllable discrimination rather than controllable generation.
@@
Figure 2: The procedure of controllable learning in Definition 2.1. Specifically, the context contains the historical interaction data, user profile, and other side information. The task target can be divided into two categories based on the source: user side and platform side, which include factors such as user mood, user privacy, and the platform’s multi-objective requirements. The task description is a representation of the task target, such as the preference weights of multi objective, natural language and so on. The controllable learner consists of the control function $h$ and the leaner $f$, taking the task description and context as input and outputting a new learner $f_{\mathcal{T}}$ to accomplish recommendation consistent of the task target.
3 Controllable Learning in Information Retrieval: Taxonomy
@@
Controllable learning plays a crucial role in modern information retrieval (IR), where learning models must dynamically adapt to various task descriptions, such as users’ specific information needs, without requiring extensive retraining. This capability ensures the delivery of personalized and relevant search results, thereby enhancing user satisfaction in IR systems. Recent advancements underscore the importance of controllability in contemporary IR platforms [19, 2, 6]. The following sections delve into the taxonomy of controllable learning in IR, exploring who controls and what is controllable in IR systems. Additionally, we also present some representative works and provide a detailed description of their control methods in terms of “what” and “who” categories. (Table 1).
@@ 3.1 What is Controllable
This section aims to clarify “what is controllable” for a controllable learner in information retrieval applications. Specifically, we explore the question: “What does the task target $\bm{s}_{\mathrm{tgt}}$ in Definition 2.1 of controllable learning could be?” The first type of task target is multi-objective control. In real-world business scenarios, both users and platforms experience shifts in interests, especially in multi-objective settings, where the emphasis on different objectives varies over time. Therefore, enabling the model to adapt to these changing objective requirements is particularly important. The second task target is user portrait control, where users can edit the context, thereby influencing the recommendation output. Some platforms provide preference questionnaires, allowing users to express personal preferences independently. Similarly, some platforms allow users to adjust their profiles and interaction history, achieving specific needs such as privacy, personalization and so on. The third task target is scenario adaptation control. Real-world platforms typically contain multiple content pages, where users’ transition data across different pages can be valuable. Additionally, in the time dimension, user behavior on some platforms is often specifically related to time segments (considered as a special scenario). Explicitly leveraging scenario information to control the recommendation system is therefore essential.
@@ 3.1.1 Multi-Objective Control
The preference (task targets) of the platform for multiple objectives (such as accuracy, diversity, and novelty) change over time. When facing a new task target, we expect a control function to directly map the learner $f$ to $f_{\mathcal{T}}$ that meets the demand based on the task description $\bm{s}_{\mathrm{{desc}}}$, where the $\bm{s}_{\mathrm{{desc}}}$ is defined as the preference for each objectives (in the explicit form of weights or natural language or the implicit form of interaction behaviors and so on). Moreover, at a finer granularity, user preferences for different categories can also be regarded as a form of multi-objective control. An example is shown in Figure 3.
Figure 3: The work flow of the Multi-Objective Control. Over time, the objectives of both the platform and users change, leading to the emergence of multi-objective requirements.
@@
ComiRec [1] strikes a balance between accuracy and diversity through hyperparameters in the aggregation module by capturing multiple user interests and integrating them into the recommendation process. UCRS [2] introduces a system where users can proactively reduce filter bubbles, controlling the recommendations they receive via hyperparameters that affect accuracy and coverage. The CMR framework [3] uses a hypernetwork to dynamically adjust preference weights, aiming to balance accuracy, diversity, etc. through platform-defined preference weights. These systems collectively focus on achieving a harmonious balance among the various objectives of recommendation quality, ensuring that users receive accurate, diverse, and relevant recommendations. PadiRec [4] treats multi-objective models under different preference weights as learning targets, considering the preference weights as conditions. Through conditional learning with a diffusion model, it captures the relationship between preference weights and the parameters of the multi-objective model. During the testing stage, by inputting the desired preference weights, the diffusion model can customize the model in real time according to these specifications, enabling control and test-time adaptation. RecLM-gen [5] utilize LLMs to follow recommendation-specific instructions. They utilize conventional recommendation models to distill recommendation capabilities into LLMs via labels. They then propose a reinforcement learning-based alignment procedure to enhance the generalization ability of LLMs, which includes category control and category proportion control.
@@ 3.1.2 User Portrait Control
Figure 4: The examples on User Portrait Control. The user portrait is reflected in various aspects such as interaction history, preference summary, and ratings. The user portrait can be modified and edited. The deletions are marked in red, and additions are marked in green.
@@
Figure 5: The work flow of the Scenario Adaptation Control. As continuous scenario switching, the control function maps the learner to a scenario-specific learner based on the task description and the scenario-specific learner is then applied to the new scenario.
@@ 3.2.1 User-Centric Control
Supervised $\beta$-VAE [15] proposes a supervised disentangled $\beta$-VAE-based recommendation model, which explicitly assigns latent dimensions to user preference attributes and aligns them with labeled features to achieve disentanglement. By incorporating classification loss for effective supervision on partially labeled data, the approach generates interpretable and controllable recommendations while maintaining high recommendation performance and enabling fine-grained user control over the output. LP [16] presents a method allowing users to explicitly control recommendation results through iterative critiques. It extends a linear embedding-based recommendation approach by integrating subjective keyphrase feedback and modulating critique strength in a multistep process. This method reduces interactions required for satisfactory recommendations and improves success rates, showcasing enhanced user-centric recommendation adaptability. TEARS [9] captures user preferences in a text-based format and leverages optimal transport to align embeddings from summaries with black-box embeddings based on user history during training. This setup allows users to edit their summaries directly, giving them control over the recommendations they receive during testing. LangPTune [17], a framework that enables explicit user control over recommendation systems through natural language profiles. By integrating end-to-end training with reinforcement learning and contrastive learning, it optimizes both recommendation accuracy and the interpretability of user preferences. This approach ensures users can directly modify their profiles to steer recommendations, enhancing transparency and personalization.
@@
UCRS [2] proposes four control commands to reduce filter bubbles. Fine-grained user control allows users to increase items associated with their specific characteristics. Coarse-grained user control helps users reduce items tied to their own profile traits. Fine-grained item control enables adjustments like increasing items from specific categories. Coarse-grained item control allows reducing items from the most frequent category in the user’s interaction history. In practical scenarios, users may not want all their behaviors to be included in the model training process. IFRQE [7] lets users decide which of their interactions should contribute to the training of the recommendation. The models are then optimized to maximize utility, which considers the trade-off between recommendation performance and respecting user preferences. LACE [6] enables users to edit concept-based profiles for personalized text recommendations. LACE represents each user as a set of human-readable concepts by retrieving items the user has interacted with. These concept-based and editable user profiles are then used to make recommendations. The model’s design enables intuitive interaction with transparent user profiles to control recommendations.

### v3

@@ 2 Foundations of Controllable Learning (CL)
In this section, we first give a formal definition of CL in terms of task requirements, and describe the procedure of CL.
@@
The concept of CL is manifested across various AI research directions, yet there is currently lack of a unified formal definition. To explicitly express the ingredients of CL and unify existing CL methods, we first define CL in terms of task specifications. In brief, CL is the ability to find a learner that can adapt to different task requirements without the need for retraining, thereby meeting the desired task targets, as detailed in Definition 2.1.
Definition 2.1 (Controllable Learning (CL)).
@@
Define a task requirement triplet $\mathcal{T}=\{\bm{s}_{\mathrm{desc}},\bm{s}_{\mathrm{ctx}},\bm{s}_{\mathrm{tgt}}\}\in\Gamma$, where $\bm{s}_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ represents the task description, $\bm{s}_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ represents the context related to the task, and $\bm{s}_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ represents the task target. Given an input space $\mathcal{X}$ and an output space $\mathcal{Y}$, for a learner $f:\mathcal{X}\rightarrow\mathcal{Y}$, controllable learning (CL) aims to find a control function $h$ that maps the learner $f$, the task description $\bm{s}_{\mathrm{desc}}\in\mathcal{T}$, and the context $\bm{s}_{\mathrm{ctx}}\in\mathcal{T}$ to a new learner $f_{\mathcal{T}}$ that fulfills the task target $\bm{s}_{\mathrm{tgt}}\in\mathcal{T}$, i.e,
@@
Unlike domain adaptation or transfer learning[83, 84, 85], CL adjusts the model adaptively in the face of new task requirements, eliminating the need for retraining for new tasks during the deployment phase. More specifically, in the context of a controllable learner during the testing phase, adjustments based on the task description can be categorized as follows: altering model inputs qualifies as a pre-processing method, modifying model parameters constitutes an in-processing method, and changing model outputs can be regarded as a post-processing method. However, it also presents challenges during the training phase in CL, particularly in guiding the control function to identify and adapt to various task requirements effectively.
For the task requirement in Definition 2.1, the task target $\bm{s}_{\mathrm{tgt}}$ can be considered as the ideal quantitative metric that the controllable learner aims to achieve. The task description $\bm{s}_{\mathrm{desc}}$ is the specific representation of task target $\bm{s}_{\mathrm{tgt}}$ that can be perceived by the control function $h$. Here is an example of task description $\bm{s}_{\mathrm{des}}$: Here is an example to illustrate the task target $\bm{s}_{\mathrm{tgt}}$ and the task description $\bm{s}_{\mathrm{des}}$:
Example 2.2.
In the e-commerce platform, the platform often have requirements for both accuracy and diversity in the output results, and their preference between the two may shift depending on business objectives. In such scenarios, task target $\bm{s}_{\mathrm{tgt}}$ can represent an expected performance target with respect to both accuracy and diversity (e.g., placing greater emphasis on diversity). Following this scenario, $\bm{s}_{\mathrm{desc}}$ can be represented either as a vector containing the weights of accuracy and diversity (e.g., $[0.4,0.6]$). Thus, at test time, platforms can express their preferences for diversity and accuracy through $\bm{s}_{\mathrm{desc}}$, which serves as the input to the control function $h$, thereby producing a learner $f$ that satisfies the target $\bm{s}_{\mathrm{tgt}}$.
The context $\bm{s}_{\mathrm{ctx}}$ in task requirements can consist of features like historical data and the current user’s profile, offering controllable learners more background knowledge for prediction.
@@
Under the definition, the procedure of CL could be shown as Figure 2. When the output space $\mathcal{Y}$ in Definition 2.1 is not labels or lists but instead generated content, it is referred to as controllable generation. In this survey, our primary focus lies on CL methods within information retrieval applications, which typically belong to controllable discrimination rather than controllable generation.
@@
Figure 2: The procedure of CL in Definition 2.1. Specifically, the context contains the historical interaction data, user profile, and other side information. The task target is categorized into three types: multi-objective control, user portrait control, and scenario adaptation control. The rationale behind this classification is detailed in Section 3.1. The task description is a representation of the task target, such as the preference weights of multi objective, natural language and so on. The controllable learner consists of the control function $h$ and the learner $f$, taking the task description and context as input and outputting a new learner $f_{\mathcal{T}}$ to accomplish recommendation consistent of the task target.
2.1 Comparison with Other Topics
Fairness, privacy, and interpretability are key aspects of trustworthy machine learning. Fairness ensures that model predictions do not exhibit bias against protected groups; privacy focuses on protecting sensitive user data from exposure; interpretability provides transparency into model behavior, enabling understanding and trust. Controllable learning (CL), by contrast, emphasizes the ability of a model to adapt to dynamic task requirements at test time without retraining, guided by a task requirement triplet consisting of task description, context, and target. While fairness and privacy are often enforced through training-time constraints, interpretability facilitates human understanding and is foundational for enabling test-time control. Within the CL framework, fairness can be formulated as a target in multi-objective control, privacy can be addressed via user portrait control by modifying context inputs, and interpretability can support interactive control through editable task descriptions in natural language or concept-based formats. Therefore, CL serves as a crucial mechanism for aligning model behavior with diverse and evolving user or platform objectives, bridging multiple dimensions of trustworthy machine learning.
3 Controllable Learning (CL) in Information Retrieval: Taxonomy
@@
CL plays a crucial role in modern information retrieval (IR), where learning models must dynamically adapt to various task descriptions, such as users’ specific information needs, without requiring extensive retraining. This capability ensures the delivery of personalized and relevant search results, thereby enhancing user satisfaction in IR systems. Recent advancements underscore the importance of controllability in contemporary IR platforms [22, 2, 6]. The following sections delve into the taxonomy of CL in IR, exploring who controls and what is controllable in IR systems. Additionally, we also present some representative works and provide a detailed description of their control methods in terms of “what” and “who” categories. (Table 1).
3.1 Why This Taxonomy
The rationale behind the taxonomy is as follows. From a practical standpoint, it is natural to ask: what do we aim to control in IR? In essence, the purpose of control can be broadly categorized into three types. First, we control for shifting interests to quickly satisfy the diverse and evolving needs of both users and platforms—this is denoted as multi-objective control. Second, we control for privacy protection by allowing users to edit their personal profiles, referred to as user-portrait control. Third, we control for environmental adaptation, enabling models to efficiently adjust to varying scenarios through explicit scenario descriptions—this is termed scenario adaptation control. In addition to the question of “what to control”, we also consider “who controls”. Both users and platforms have control needs, but as initiators and recipients of IR systems respectively, their control intents diverge. Hence, we distinguish their roles and objectives accordingly in our taxonomy. Finally, we delve into the technical dimension to categorize and summarize common approaches to CL, which are captured by the axes of “How” (i.e., control techniques) in Section  4.2 and “Where” (i.e., the stage in the pipeline where control is applied) in Section 4.1.
Importantly, these practical needs directly correspond to the formal components in our definition of CL. We unify the CL process into a framework as illustrated in Figure 2, following our definition of CL. As shown in Definition 2.1, the execution of CL typically requires a clearly specified task requirement. We define the task requirement triplet as follows: $\mathcal{T}={\bm{s}_{\mathrm{desc}},\bm{s}_{\mathrm{ctx}},\bm{s}_{\mathrm{tgt}}}\in\Gamma$, where $\bm{s}_{\mathrm{desc}}\in\mathcal{D}_{\mathrm{desc}}$ denotes the task description, $\bm{s}_{\mathrm{ctx}}\in\mathcal{D}_{\mathrm{ctx}}$ denotes the task-related context, and $\bm{s}_{\mathrm{tgt}}\in\mathcal{D}_{\mathrm{tgt}}$ denotes the task target. For the task target $\bm{s}_{\mathrm{tgt}}\in\mathcal{T}$, we explore the question of “control for what” in the IR scenario, i.e., “what” the target $\bm{s}_{\mathrm{tgt}}$ could be. However, the two parties in an IR system—users and platforms—often focus on different aspects of $\bm{s}_{\mathrm{tgt}}$, leading to divergent task formulations (i.e., $\bm{s}_{\mathrm{des}}$). Thus, we further investigate “who” (user or platform) is responsible for specifying the task description under different target settings, as this sheds light on their respective controllability intentions. Finally, to support future research, we categorize common techniques for CL (denoted as “How” ), and map their functional roles to the corresponding modules of the CL framework (denoted as “Where”) to facilitate structured understanding and community advancement.
@@ 3.2 What is Controllable
This section aims to clarify “what is controllable” for a controllable learner in information retrieval applications. Specifically, we explore the question: “What does the task target $\bm{s}_{\mathrm{tgt}}$ in Definition 2.1 of CL could be?” The first type of task target is multi-objective control. In real-world business scenarios, both users and platforms experience shifts in interests, especially in multi-objective settings, where the emphasis on different objectives varies over time. Therefore, enabling the model to adapt to these changing objective requirements is particularly important. The second task target is user portrait control, where users can edit the context, thereby influencing the recommendation output. Some platforms provide preference questionnaires, allowing users to express personal preferences independently. Similarly, some platforms allow users to adjust their profiles and interaction history, achieving specific needs such as privacy, personalization and so on. The third task target is scenario adaptation control. Real-world platforms typically contain multiple content pages, where users’ transition data across different pages can be valuable. Additionally, in the time dimension, user behavior on some platforms is often specifically related to time segments (considered as a special scenario). Explicitly leveraging scenario information to control the recommendation system is therefore essential.
@@ 3.2.1 Multi-Objective Control
The preference (task targets) of the platform for multiple objectives (such as accuracy, diversity, and novelty) change over time. When facing a new task target, we expect a control function to directly map the learner $f$ to $f_{\mathcal{T}}$ that meets the demand based on the task description $\bm{s}_{\mathrm{{desc}}}$, where the $\bm{s}_{\mathrm{{desc}}}$ is defined as the preference for each objectives (in the explicit form of weights or natural language or the implicit form of interaction behaviors and so on). Moreover, at a finer granularity, user preferences for different categories can also be regarded as a form of multi-objective control. Figure 3 illustrates an example of multi-objective control. Assuming the task target includes objectives related to accuracy and diversity, and the task description is represented as a two-dimensional vector indicating preferences for these objectives. From the platform’s perspective, there is a strong emphasis on accuracy in a given period, leading to homogeneous recommendations (as shown by similar circles being recommended). Suppose the platform adjusts its business metrics to increase diversity in the output results, indicating a change in the task target. According to the definition of CL, we only need to modify the task description (i.e., increase the value of the diversity dimension in the two-dimensional vector). The control function can then directly map the learner to meet the new task target, instead of retraining the model based on the new task target. On the user side, if we consider preferences for different types of items as different objectives, a similar analysis can be applied as in the platform-side case above.
Figure 3: Illustration of the need for Multi-Objective Control. Users may have a temporary preference during the test stage (e.g., shifting from “Love” and “Suspense” to “Fiction” after a temporary description). The platform may instead focus more on the diversity of the outputs rather than accuracy during the test stage. This figure depicts the dynamic nature of shifting objectives for both the platform and users during test stage, underscoring the critical role of multi-objective control in recommender systems.
@@
ComiRec [1] strikes a balance between accuracy and diversity through hyperparameters in the aggregation module by capturing multiple user interests and integrating them into the recommendation process. UCRS [2] introduces a system where users can proactively reduce filter bubbles, controlling the recommendations they receive via hyperparameters that affect accuracy and coverage. Although the above two algorithms primarily focus on balancing multiple objectives, they also enable the possibility of dynamic control. For example, ComiRec introduces a controllable hyperparameter $\lambda$ in the multi-objective aggregation module to adjust diversity, and UCRS provides control coefficients $\alpha$ and $\beta$ to regulate accuracy, isolation, and diversity. These adjustments on hyperparameters do not trigger model retraining, thus providing a foundation for dynamic control during the test stage. The CMR framework [3] is specifically designed to address the dynamic nature of preferences across multiple objectives. In this framework, a hypernetwork takes dynamic preference weights as input to adjust the last few layers of the recommendation model, enabling dynamic control over objectives such as accuracy and diversity. These systems collectively focus on achieving a harmonious balance among the various objectives of recommendation quality, ensuring that users receive accurate, diverse, and relevant recommendations. PadiRec [4] treats multi-objective models under different preference weights as learning targets, considering the preference weights as conditions. Through conditional learning with a diffusion model, it captures the relationship between preference weights and the parameters of the multi-objective model. During the testing stage, by inputting the desired preference weights, the diffusion model can customize the model in real time according to these specifications, enabling control and test-time adaptation. RecLM-gen [5] utilize LLMs to follow recommendation-specific instructions. They utilize conventional recommendation models to distill recommendation capabilities into LLMs via labels. They then propose a reinforcement learning-based alignment procedure to enhance the generalization ability of LLMs, which includes category control and category proportion control.
@@ 3.2.2 User Portrait Control
Figure 4: The Examples on User Portrait Control. The preference summary represents an aggregation of the user’s interaction history. The list on the right shows the output items along with their ratings. Deletions are marked in red, and additions are marked in green. This figure illustrates three forms of user modification, demonstrating that user portrait control can be achieved by modifying and editing the interaction history, preference summary, and ratings.
@@
Figure 5: The Workflow of Scenario Adaptation Control. Unlike traditional settings where a learner is fixed to a specific scenario, Scenario Adaptation Control aims to address the challenge of learner adaptation under continuous scenario switching during the test stage. When a new scenario arises, the control function maps the general learner to a scenario-specific learner based on the corresponding task description. This allows the scenario-specific learner to be applied to the new context without the need for retraining.
@@ 3.3.1 User-Centric Control
Supervised $\beta$-VAE [15] proposes a supervised disentangled $\beta$-VAE-based recommendation model, which explicitly assigns latent dimensions to user preference attributes and aligns them with labeled features to achieve disentanglement. By incorporating classification loss for effective supervision on partially labeled data, the approach generates interpretable and controllable recommendations while maintaining high recommendation performance and enabling fine-grained user control over the output. LP [16] presents a method allowing users to explicitly control recommendation results through iterative critiques. It extends a linear embedding-based recommendation approach by integrating subjective keyphrase feedback and modulating critique strength in a multistep process. This method reduces interactions required for satisfactory recommendations and improves success rates, showcasing enhanced user-centric recommendation adaptability. TEARS [9] captures user preferences in a text-based format and leverages optimal transport to align embeddings from summaries with black-box embeddings based on user history during training. This setup allows users to edit their summaries directly, giving them control over the recommendations they receive during testing. LangPTune [17], a framework that enables explicit user control over recommendation systems through natural language profiles. By integrating end-to-end training with reinforcement learning and contrastive learning, it optimizes both recommendation accuracy and the interpretability of user preferences. This approach ensures users can directly modify their profiles to steer recommendations, enhancing transparency and personalization. GOMMIR [18] incorporates user language as a reward signal into reinforcement learning optimization, enabling explicit user-side controllability through goal-oriented rewards and further optimizing recommendation performance. Promptriever [19] introduces the first retrieval model capable of being prompted like a language model. It not only achieves strong performance on standard retrieval tasks but also follows instructions, enabling explicit user control.
@@
UCRS [2] proposes four control commands to reduce filter bubbles. Fine-grained user control allows users to increase items associated with their specific characteristics. Coarse-grained user control helps users reduce items tied to their own profile traits. Fine-grained item control enables adjustments like increasing items from specific categories. Coarse-grained item control allows reducing items from the most frequent category in the user’s interaction history. In practical scenarios, users may not want all their behaviors to be included in the model training process. IFRQE [7] lets users decide which of their interactions should contribute to the training of the recommendation. The models are then optimized to maximize utility, which considers the trade-off between recommendation performance and respecting user preferences. LACE [6] enables users to edit concept-based profiles for personalized text recommendations. LACE represents each user as a set of human-readable concepts by retrieving items the user has interacted with. These concept-based and editable user profiles are then used to make recommendations. The model’s design enables intuitive interaction with transparent user profiles to control recommendations. InstructAgent [20] introduces a new “User-Agent-Platform” paradigm and proposes an instruction-aware agent, InstructAgent, which acts as an intermediary between users and the recommendation system. It addresses issues in traditional recommendation systems such as compromised user interests and insufficient personalization, thereby providing a new paradigm for user control.